A traveler needs a rental car for a day. He shops around to find a car rental company with the lowest overall price. He finds a national car rental company that charges a $24 daily fee plus $0.80 per mile. He also finds a local car rental company that charges a $18 daily fee, plus $0.90 per mile. Under what conditions would it be less expensive to rent from the national car rental company?
Answer #1

Let the distance travelled is P miles

Total bill generated through National car rental

Amount = 24 + 0.80P

Total bill generated through Local car rental

Amount = 18 + 0.90P

Under conditions of indifference, both car rentals should be of same amount for P miles

24 + 0.80P = 18 + 0.90P

6 = 0.10P

p = 60 miles

Hence, when the miles travelled is more than P=60miles, it would be less expensive to travel through National car rental company.